OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sine of five pi divided by twelve A: one half times the square root of quantity two plus square root of three B: the square root of quantity two plus square root of three C: one half times the square root of quantity two minus square root of three D: the square root of quantity two minus square root of three

zepdrix (zepdrix):

\[\large\rm \sin\left(\frac{5\pi}{12}\right)=\sin\left(\frac{\color{orangered}{5\pi/6}}{2}\right)\]

zepdrix (zepdrix):

Recall your Half-Angle Identity:\[\large\rm \sin\left(\frac{\color{orangered}{\theta}}{2}\right)=\pm\sqrt{\frac{1-\cos\color{orangered}{\theta}}{2}}\]

OpenStudy (anonymous):

yep i remeber it but i am stuck on this one problem and have been stuck for hours need to get it over with

OpenStudy (anonymous):

where you see a \(\theta\) put \(\frac{5\pi}{6}\) that is all

OpenStudy (anonymous):

do you know \[\cos(\frac{5\pi}{6})\]?

OpenStudy (anonymous):

"no" is a fine answer, just asking

OpenStudy (anonymous):

 right?